OpenStudy (anonymous):

What is the next number in the pattern below? 3, 5, 9, 15, 23, ___ 31 33 35 37

OpenStudy (wil476003):

It would be 33 because 3+2=5 5+4=9 9+6=15 15+8=23 23+10=33 so 33 is the answer I hope my this makes sense to you the work I did otherwise I can explain it more